OpenStudy (anonymous):

PLEASE HELP!! FAN AND MEDAL! Solve. 5 = –(–z + 3) A. –8 B. 8 C. –2 D. 2

OpenStudy (lynfran):

5=-(-z+3) distribute the negative sign 5=z-3 add 3 to both sides we get 5+3=z-3+3 5+3=z